# VALUE: Value-Aware Large Language Model for Query Rewriting via Weighted Trie in Sponsored Search

### v1

@@ 3.2.1. Definition and Construction of Weighted Trie
Our Weighted Trie is an renovation of the traditional trie data structure. Each node in the Weighted Trie, referred to as a WeightedTrieNode, contains additional attributes to store weight-related information. The primary attributes of a WeightedTrieNode are:
  • children: A dictionary where the keys are integers representing tokens, and the values are child nodes of type WeightedTrieNode.
  • mean: A floating-point number representing the average weight of the node’s children.
  • max: A floating-point number representing the maximum weight among the node’s children.
@@
Construction of Weighted Trie: the construction of the Weighted Trie involves initializing the root node, inserting bidwords with their corresponding values, and updating the weights of the nodes. The pseudocode in Algorithm 1 demonstrates the construction process.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.2.2. Weighted Trie Momentum Update
Momentum updates are essential due to the frequent changes in eCPM and the evolving bidword space. The process is delineated as follows:
@@
(1) $V_{\text{new}}=\max(eCPM_{\text{new}},V_{\text{old}})$
(2) $W_{\text{new}}=\alpha\cdot eCPM_{\text{new}}\ +\ \beta\cdot W_{\text{old}}$
where $\alpha$ and $\beta$ are hyperparameters that control the update rate, with the constraint $\alpha+\beta=1$. V is the max attribute of the node and W is the mean attribute of the node. If the node is absent, it is initialized as:
(3) $W_{\text{new}}=eCPM_{\text{new}},\ V_{\text{new}}=eCPM_{\text{new}}$
@@ 3.3. Collaborative Inference
As shown in Figure 2, when generating the $t$-th token, we insert the query into the prompt template and concatenate it with the bidword prefix to form the input of LLM to get the output probabilities. We use generated bidword prefix to retrieve all child nodes from the trie. The value of these child nodes is computed using the following formula:
(4) $\mathcal{V}_{k}=\alpha\cdot\text{mean}_{k}+\beta\cdot\text{max}_{k}$
@@
Here, $\alpha$ and $\beta$ are tunable hyperparameters, both set to 0.5, allowing us to prioritize paths with higher average values or to explore bidwords with the maximum potential value, depending on our specific needs.
@@
We then apply the softmax function to these values to obtain the normalized value $\hat{\mathcal{V}}_{k}$. Next, we mask out all probabilities in the LLM-generated distribution that are not children of the current node. For the remaining probabilities, we adjust them by weighting with the normalized value $\hat{\mathcal{V}}_{k}$ and a depth-dependent factor $\theta$. The adjusted probability for each token $k$ is given by:
(5) $p(k\mid x_{<t})=p_{\text{LLM}}(k\mid x_{<t})\cdot(1+\mathcal{V}_{k}\cdot\theta)$
where $\theta$ is designed to vary with the depth of the trie, playing a crucial role in balancing relevance and value in our model’s output. At shallower levels of the trie, $\theta$ is set to smaller values to minimize the influence of value on the output, thereby preventing the model from favoring high-value but less relevant options. As we delve deeper into the trie, $\theta$ increases, allowing us to prioritize higher-value candidates from a pool of already highly relevant choices. Consequently, we derive a modified probability distribution $p(k\mid x_{<t})$, from which the next token is sampled as usual.
3.4. Multi-task Fine-Tuning
As no open-source LLMs are trained to perform query rewriting for the e-commerce scenario, they lack an awareness of the bidword space. Consequently, they are unable to effectively generate bidwords. To address this, we need to perform supervised fine-tuning (SFT) on these models to inject domain-specific knowledge into the model and enhancing their capability to understand and effectively rewrite e-commerce queries for improved user experience and search accuracy.
Query Rewriting Dataset: To construct our training dataset, we initially obtain query-bidword pairs from online logs. Specifically, when a user searches for a query $x$ on our e-commerce platform, the system logs the corresponding bidwords $Y=\{y^{1},y^{2},\ldots,y^{n}\}$ that are associated with the advertisements the user interacts with. From this list, we select the top-ranked bidword $y^{1}$ as the gold standard candidate to construct our initial dataset $D$ with $N$ samples:
(6) $D=\{(x_{i},y_{i}^{1})\}_{i=1}^{N}\quad\text{where}\quad x_{i}\sim p(x)$
where $p(x)$ denotes the query distribution in our search engine.
We apply a relevance filter to $D$ using a representation model fine-tuned from an open-source embedding model. The relevance score $\text{rel}(x,y)$ is computed as the cosine similarity between the embeddings of $x$ and $y$:
(7) $\text{rel}(x,y)=\frac{\text{emb}(x)\cdot\text{emb}(y)}{\|\text{emb}(x)\|\|\text{emb}(y)\|}$
We then filter the dataset $D$ to obtain $D_{r}$ ensuring that the relevance score exceeds a threshold $\tau_{\text{rel}}$:
(8) $D_{r}=\{(x_{i},y_{i})\mid(x_{i},y_{i})\in D,\,\text{rel}(x_{i},y_{i})>\tau_{\text{rel}}\}$
For head queries, the number of corresponding bidwords can be in the tens of thousands. To ensure class balance, we truncate the bidwords for each query based on their value.
Fine-tuning Tasks: We employ two types of supervised fine-tuning (SFT) tasks to enhance the model’s understanding and performance in query rewriting for e-commerce.
Task 1: <prompt + query + bidword>
This task aims to enhance the model’s spatial awareness of the bidword space and the association between queries and bidwords. Subsequently, the model will develop fundamental rewriting skills.
Task 2: <prompt + query + bidword_list>
In this task, the model is trained with lists of bidwords sorted in descending order based on their eCPM values. The bidword_list for each query is structured as follows:
[…]
@@ 3.5. LLM Post-training Alignment with Weighted Trie
Although conventional RLHF approaches can effectively integrate with our framework, we propose a more tailored solution. Our Weighted DPO (WDPO) approach directly incorporates the economic value of bidwords into the alignment process, leveraging the eCPM values for bidwords. It is important to note that WDPO is optional as the performance improvements it offers are not as significant relative to computational cost when compared to SFT and the varying nature of bidwords. Therefore, WDPO can be considered when resources permit, providing an additional layer of refinement to the training process.
@@ 3.6. Online Depolyment
In the context of deploying LLMs for online serving, the balance between computational efficiency and latency is paramount. Our approach leverages the Qwen2-7B model for offline inference, targeting head queries that constitute 50% of the page views (PV) in our system. This offline process generates the top 500 results per query, which are subsequently cached for rapid access. For mid-tail and long-tail queries that do not benefit from the cached results, we employ the Qwen2-1.5B model in an online serving capacity. This model is optimized to deliver a latency of 50 milliseconds, efficiently producing the top 50 results per query. In the future, we plan to integrate user behavior sequence data into our online serving framework to enhance personalized recommendation systems and gradually remove the offline part.
@@ 4.1. Datasets
[…]
@@ 4.2. Evaluation Metrics
[…]
@@ 4.3. Implementation Details
Multi-task Fine-tuning: We fine-tuned the Qwen2-7B model using a learning rate of 1e-5 and a cosine learning rate scheduler. The process was optimized with AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a weight decay of 0.001. Gradient accumulation was set to 4. We used a batch size of 128 and applied the DeepSpeed ZeRO Stage 2 parallelism strategy with bfloat16 precision. The training was performed on 32 NVIDIA H20 GPUs for a total of 12 hours.
@@ 4.4. Offline Experiments
Model hitrate@50 hitrate@500 Spearman’s $\rho$ Relevance OOVR eCPM
OnlineModel 59.93% 85.41% 0.02 68.15% 0% 8676
DeepNB 44.16% 79.82% 0.13 46.26% 0% 12852
BEQUE 39.26% 62.85% 0.08 74.37% 38.41% 7582
raw Qwen2-7B 12.88% 18.37% -0.03 59.81% 68.63% 2937
SFT Qwen2-7B 38.18% 63.71% 0.02 76.33% 34.85% 7439
SFT+Trie 59.11% 86.38% -0.01 75.31% 0% 7270
SFT+Trie+DPO 38.61% 62.19% 0.05 69.48% 0% 28391
SFT+WeightedTrie(VALUE) 60.37% 91.52% 0.46 74.55% 0% 59803
SFT+WeightedTrie+DPO 50.35% 77.73% 0.23 70.18% 0% 38301
SFT+WeightedTrie+WDPO(VALUE+WDPO) 62.67% 91.91% 0.56 73.49% 0% 63775
In this section, we present the results of offline experiments conducted to evaluate the performance of our proposed models against several baseline models. The comparative performance analysis is summarized in Table 1. We compared our models with the following baselines: first, the Online Model, which represents the current state-of-the-art multi-channel recall system deployed online; second, the DeepNB model (Chen et al., 2020), which employs a vector-based approach focusing solely on value; and finally, the BEQUE model (Peng et al., 2023), which utilizes a LLM fine-tuned with multi-task SFT and PRO alignment, representing the latest query rewriting approach. The performance metrics used for evaluation include hitrate, eCPM, Spearman’s $\rho$, Relevance, and OOVR(Out-Of-Vocabulary Rate. To a certain extent, it is also equivalent to the hallucination rate of LLMs.). These metrics provide a comprehensive assessment of the models’ effectiveness in terms of query rewriting, relevance, and OOVR.
@@ 4.4.1. Comprehensive Evaluation of the VALUE Framework
The VALUE framework we propose is designed to optimize online RPM revenue by utilizing offline eCPM values as a reference for learning effectiveness, rather than as direct revenue indicators. Our subsequent online experiments have demonstrated the validity and rationality of this modeling approach.The results presented in Table 1 clearly illustrate the significant effectiveness of our VALUE framework. Notably, the eCPM metric shows a remarkable improvement, serving as a strong indicator of offline learning success. Compared to the model trained with Supervised Fine-Tuning (SFT) on the same dataset, our VALUE model achieves a 722.6% increase in eCPM. Furthermore, when compared to the current online service model, our framework exhibits a 589.3% enhancement in eCPM. Importantly, the relevance metric for our VALUE model does not decline when compared to the SFT model; in fact, it demonstrates an improvement of 6.4 percentage points over the online service model. This observation substantiates our earlier hypothesis that relevance and value are decoupled in this domain, allowing us to reorder the generation sequence of bidwords to enhance both metrics simultaneously without trade-offs.
Compared to the DeepNB model, which focuses on efficiency-oriented rewriting using click data to maximize value, our VALUE framework demonstrates superior performance across multiple dimensions. While the DeepNB model effectively maximizes value, it suffers from poor relevance metrics. In contrast, our VALUE framework not only maintains high relevance but also significantly outperforms DeepNB in terms of value attributes. Specifically, the eCPM metric for our VALUE model is substantially higher than that of DeepNB. When compared to the BEQUE model, our framework achieves comparable relevance metrics but excels in generating high-value bidwords. This is evidenced by the eCPM metric, where VALUE achieves a value of 59803, far surpassing BEQUE’s 7582. Additionally, our VALUE framework demonstrates a superior hit rate, with hitrate@500 reaching 91.52%, compared to BEQUE’s 62.85%. This significant improvement in hit rate underscores the effectiveness of our approach in generating bidwords that are not only relevant but also highly likely to be clicked by users. It is noteworthy that the BEQUE method suffers from a high Out-Of-Vocabulary Rate (OOVR) of 38.41%, which adversely affects its hit rate.
Moreover, our VALUE framework outperforms the online service model across all evaluated metrics, further validating the efficacy of our approach. The comprehensive improvements in eCPM, relevance, and other metrics underscore the robustness and effectiveness of our method in generating high-value and contextually relevant bidwords for e-commerce query rewriting tasks.
@@
Figure 3. Trade-off between RPM and PV rele through $\theta$.
[…]
@@
$\theta_{1}$ 86.38% 75.31%
$\theta_{2}$ 89.65% 75.04%
$\theta_{3}$ 91.52% 74.55%
$\theta_{4}$ 88.63% 69.85%
$\theta_{5}$ 73.71% 60.53%
4.4.2. Enhanced Performance in Fine-Grained Value Alignment
Compared to traditional alignment methods such as DPO and Preference Ranking Optimization (PRO) used in the BEQUE model, our VALUE framework demonstrates a significant lead in the Spearman rank correlation metric. This indicates that conventional value alignment methods struggle with fine-grained value alignment tasks. Specifically, the Spearman’s $\rho$ for our VALUE model is 0.46, which is substantially higher than the 0.05 achieved by DPO model and 0.08 by BEQUE.
Additionally, our framework seamlessly integrates with conventional model alignment methods. When we combined our proposed WDPO with the VALUE framework, we achieved the highest Spearman’s $\rho$ of 0.56.
DPO suffers from significant overfitting issues, which severely impact its performance in terms of hit rate and relevance. In our experiments, the SFT + DPO + Trie model exhibited a hitrate@500 of only 62.19% and a relevance score of 69.48%, both of which are considerably lower than those achieved by our VALUE framework. DPO faces inherent challenges when dealing with positive and negative samples that are lexically very similar. The overfitting problem in DPO arises because the method tends to excessively optimize for the specific preference pairs seen during training, leading to a decline in generalization capability. DPO enhances the model’s belief in specific events. Therefore, when we use DPO in conjunction with our VALUE framework, the improvement in eCPM is not significant. Specifically, the eCPM for the VALUE + DPO model is 38301, which is only marginally higher than the 28391 achieved by the SFT + DPO + WeightedTrie model.
Furthermore, the improvement in the Spearman rank correlation metric with DPO is not significant. The SFT + DPO + Trie model achieved a Spearman’s $\rho$ of only 0.05, which is marginally better than the baseline but far from satisfactory. This limited improvement underscores the inherent limitations of DPO in handling fine-grained value alignment tasks.

### v2

@@ 3.1.1. Definition and Construction of Weighted Trie
Weighted Trie is an renovation of the traditional trie data structure. Each node in the Weighted Trie, referred to as a WeightedTrieNode, contains additional attributes to store weight-related information. The primary attributes of a WeightedTrieNode are:
  • children: A dictionary where the keys are integers representing tokens in LLM’s vocabulary, and the values are child nodes of type WeightedTrieNode.
  • mean_v: A floating-point number representing the average eCPM value of the node’s children, where eCPM (effective Cost Per Mille) is the bidword value calculated offline.
  • max_v: A floating-point number representing the maximum eCPM value among the node’s children.
@@
The construction of the Weighted Trie involves initializing the root node, inserting bidwords with their corresponding values, and updating the weights of the nodes. The pseudocode in Algorithm 1 demonstrates the construction process.
@@
Input: List of pairs $(\textit{bidword},\textit{ecpm})$
@@
 Initialize weighted trie: $\mathit{weighted\_trie}$
for each $(\textit{bidword},\textit{ecpm})$ in list do
  $\mathit{token\_ids}\leftarrow\text{tokenizer}(\textit{bidword})$
  Insert $\mathit{token\_ids}$ into the trie:
  for each $\mathit{token\_id}$ in $\mathit{token\_ids}$ do
   Traverse or create node for $\mathit{token\_id}$
@@
  Set $\mathit{leaf\_node.mean\_v}\leftarrow\textit{ecpm}$
  Set $\mathit{leaf\_node.max\_v}\leftarrow\textit{ecpm}$
@@
  $\mathit{node.mean\_v}\leftarrow\frac{1}{N}\sum_{\mathit{child}\in\mathit{children}}\mathit{child.mean\_v}$
  $\mathit{node.max\_v}\leftarrow\max_{\mathit{child}\in\mathit{children}}\mathit{child.max\_v}$
@@ 3.1.2. Weighted Trie Momentum Update
Momentum updates are essential due to the frequent changes in bidding prices and the evolving bidword space. The process is delineated as follows:
@@
(1) $Max\_V_{\text{new}}=\max(eCPM_{\text{new}},Max\_V_{\text{old}})$
(2) $Mean\_V_{\text{new}}=\alpha\cdot eCPM_{\text{new}}\ +\ (1-\alpha)\cdot Mean\_V_{\text{old}}$
where $\alpha$ is the hyperparameter that control the update rate. If the node is absent, it is initialized as:
(3) $Max\_V_{\text{new}}=eCPM_{\text{new}},\ Mean\_V_{\text{new}}=eCPM_{\text{new}}$
@@ 3.2. Collaborative Inference
As shown in Figure 2, when generating the $t$-th token, we insert the query into the prompt template and concatenate it with the bidword prefix(i.e. the first t-1 tokens) to form the input of LLM to get the output probabilities. We use generated bidword prefix to retrieve all child nodes from the trie. The value of these child nodes is computed using the following formula:
(4) $\mathcal{V}_{k}=\beta\cdot Max\_V_{k}+(1-\beta)\cdot Mean\_V_{k}$
@@
where $\beta$ is the tunable hyperparameter set to 0.5, allowing us to prioritize paths with higher average values or to explore bidwords with the maximum potential value, depending on our specific needs.
@@
We then apply the softmax function to these values to obtain the normalized value $\hat{\mathcal{V}}_{k}$. Next, we mask out all probabilities in the LLM-generated distribution that are not children of the previous generated token in the Weighted Trie. For the remaining probabilities, we adjust them by weighting with the normalized value $\hat{\mathcal{V}}_{k}$ and a depth-dependent factor $\theta$. The adjusted probability for each token $k$ is given by:
(5) $p_{WT}(k\mid x_{<t})=p_{{LLM}}(k\mid x_{<t})\cdot(1+\hat{\mathcal{V}}_{k}\cdot\theta_{t})$
The fusion value $p_{WT}(k\mid x_{<t})$ is an affine transformation of $p_{{LLM}}(k\mid x_{<t})$ and $\hat{\mathcal{V}}_{k}$, selected through a series of tests and comparisons with additive fusion, multiplicative fusion, etc. where $\theta_{t}\in\Theta$ is designed to vary with the depth $t$ of the trie, playing a crucial role in balancing relevance and value in our model’s output. Specifically, $\Theta$ assigns lower value weight to initial generated tokens and higher weight to later generated tokens. This design stems from two key observations: (1) Initial tokens primarily establish semantic relevance. Excessive value focus at this stage risks compromising query intent alignment. (2) Subsequent tokens refine bidword specificity. Once core semantics are anchored, increased value weighting allows precise optimization within the relevant search space. Thus, $\Theta$ progressively shifts focus from relevance grounding to value maximization across the generation trajectory. Consequently, we derive a modified probability distribution $p_{WT}(k\mid x_{<t})$, from which the next token is sampled as usual.
@@ 3.3. LLM Post-training Alignment with Weighted Trie
Although the Weighted Trie offers a lightweight mechanism for balancing relevance and commercial value of rewriting results, we seek to further enhance the LLM’s generation probability for high-value bidwords. This is critical because, in scenarios where a query has numerous relevant bidwords, high-value candidates meeting the relevance threshold might be excluded from the rewrite set due to low generation probabilities, thereby limiting the effectiveness of the Weighted Trie. However, prevailing value alignment techniques, exemplified by DPO, often excessively prioritize bidwords value ranking. This over-emphasis can lead to model overfitting or catastrophic forgetting of the semantic knowledge acquired during SFT.
To address these challenges, we propose Weighted Direct Preference Optimization (WDPO). WDPO directly incorporates bidwords value into DPO, steering the model’s attention toward bidword pairs exhibiting substantial value disparities. This approach effectively mitigates the negative impact of the value alignment phase on rewriting relevance.
@@ 3.4. Online Depolyment
We constructed a WeightedTrie using a dataset of 30 million bidwords. In real-world business scenarios, the growth of the bidwords is typically gradual and has negligible impact on query processing performance. As an alternative, a pruning strategy based on key metrics such as eCPM or historical number of clicks can be employed to maintain the bidword count within predefined performance bounds.
In the context of deploying LLMs for online serving, the balance between computational efficiency and latency is paramount. Our approach leverages the Qwen2-7B(Team, 2024) model for offline inference, targeting head queries that constitute 85% of the page views (PV) in our system. This offline process generates the top 500 bidwords per query, which are subsequently cached for rapid access. For mid-tail and long-tail queries that do not benefit from the cached results, we employ the Qwen2-1.5B model in an online serving capacity. This model is optimized to deliver a latency of 50 milliseconds, efficiently producing the top 50 bidwords per query. In the future, we plan to integrate user behavior sequence data into our online serving framework to enhance personalized recommendation systems and gradually remove the offline part.
@@ 4.1. Datasets
Training Dataset: For multi-task fine-tuning, we extracted 150 billion records from 30 days of online logs, followed by several rounds of data cleaning. Initially, we employed manual observation and regular expressions to filter out the majority of noisy queries. In the subsequent screening phase, we calculated the click-through rate (CTR) based on page views (PV) and clicks, retaining only those $\langle\texttt{query},\texttt{bidword}\rangle$ pairs associated with purchase behavior that meet the click-through rate requirements. We then applied a relevance model for further filtering. This relevance model is a fine-tuned version of BAAI/bge-large-zh-v1.5(Xiao et al., 2024), optimized for e-commerce scenarios. We truncated the number of bidwords for each query, retaining a maximum of 50 bidwords per query. After third round of filtering, we were left with 110 million records.
@@ 4.2. Evaluation Metrics
    (10) $\text{hitrate@}K=\frac{\sum_{q\in\mathcal{Q}}|\mathcal{R}_{q}^{K}\cap\mathcal{C}_{q}|}{\sum_{q\in\mathcal{Q}}|\mathcal{C}_{q}|}$
@@ 4.3. Implementation Details
Supervised Fine-tuning: We fine-tuned the Qwen2-7B model using a learning rate of 1e-5 and a cosine learning rate scheduler. The process was optimized with AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a weight decay of 0.001. Gradient accumulation was set to 4. We used a batch size of 128 and applied the DeepSpeed ZeRO Stage 2 parallelism strategy with bfloat16 precision. The training was performed on 32 NVIDIA H20 GPUs for a total of 12 hours.
@@
Construction and Update of Weighted Trie: We built a Weighted Trie containing 30 million bidwords with an average length of 7.9 characters, which occupies approximately 2.7 GB of memory after serialized. Single-threaded construction finishes in 56 minutes on a 32-core CPU machine, but multi-threaded optimization reduces this time to roughly 20 minutes. We used a message queue for streaming updates to the Weighted Trie, achieved minute-level update latency.
@@ 4.4. Offline Experiments
Experiment Number Model hitrate@50 hitrate@500 Spearman’s $\rho$ Relevance OOVR eCPM
Exp.1 OnlineModel 59.93% 85.41% 0.02 68.15% 0% 8676
Exp.2 DeepNB 44.16% 79.82% 0.13 46.26% 0% 12852
Exp.3 BEQUE 39.26% 62.85% 0.08 74.37% 38.41% 7582
Exp.4 BEQUE+Trie 42.77% 65.36% 0.04 72.96% 0% 7225
Exp.5 BEQUE+WeightedTrie 53.94% 79.20% 0.26 70.73% 0% 19275
Exp.6 raw Qwen2-7B 12.88% 18.37% -0.03 59.81% 68.63% 2937
Exp.7 SFT Qwen2-7B 38.18% 63.71% 0.02 76.33% 34.85% 7439
Exp.8 SFT+Trie 59.11% 86.38% -0.01 75.31% 0% 7270
Exp.9 SFT+DPO+Trie 38.61% 62.19% 0.05 69.48% 0% 28391
Exp.10 SFT+WeightedTrie 60.37% 91.52% 0.46 74.55% 0% 59803
Exp.11 SFT+DPO+WeightedTrie 50.35% 77.73% 0.23 70.18% 0% 38301
Exp.12 SFT+WDPO+WeightedTrie(VALUE) 62.67% 91.91% 0.56 73.49% 0% 63775
In this section, we present the results of offline experiments conducted to evaluate the performance of our proposed models against several baseline models. The comparative performance analysis is summarized in Table 1. We compared our models with the following baselines: first, the Online Model, which represents the current state-of-the-art multi-channel recall system deployed online; second, the DeepNB model (Chen et al., 2020), which employs a vector-based approach focusing solely on bidword value; and finally, the BEQUE model (Peng et al., 2023), which utilizes a LLM fine-tuned with multi-task SFT and PRO alignment(Song et al., 2024), representing the generative query rewriting approach. The performance metrics used for evaluation include hitrate, eCPM, Spearman’s $\rho$, Relevance, and OOVR(Out-Of-Vocabulary Rate. To a certain extent, it is also equivalent to the hallucination rate of LLMs.). These metrics provide a comprehensive assessment of the models’ effectiveness in terms of query rewriting, relevance, and OOVR.
@@ 4.4.1. Comprehensive Evaluation of the VALUE Framework
As evidenced in Table 1, VALUE (Exp.12) demonstrates superior performance across all evaluation metrics, underscoring the robustness and effectiveness of our proposed method in generating high-value and contextually relevant bidwords for e-commerce query rewriting. DeepNB (Exp.2) leverages click logs for efficiency-oriented rewriting to maximize value, yet its rewrites suffer from compromised relevance to the original search queries. Conversely, the generative approach BEQUE (Exp.3) capitalizes on the semantic comprehension capabilities of LLMs, achieving high rewrite relevance, but exhibits poor value awareness regarding bidwords. In contrast to both methods, VALUE significantly enhances the eCPM and hitrate of the rewritten queries while maintaining high relevance. This indicates that VALUE effectively synthesizes the complex semantic understanding derived from LLMs with value awareness enabled by the Weighted Trie. Notably, BEQUE exhibits a remarkably high Out-of-Vocabulary Word Rate (OOVR) of 38.41%, and integrating the Trie-based inference structure into BEQUE (Exp.4) significantly enhances hitrate while ensuring output authenticity. Furthermore, compared to the mainstream generative method ”SFT+Trie+DPO” (Exp.9), VALUE achieves improvements across all evaluation metrics. This outcome suggests that our proposed reward-guided inference framework mitigates potential overfitting inherent in the value alignment stage.
4.4.2. Ablation Study of the Weighted Trie.
Our proposed Weighted Trie guides LLMs during the generation phase to produce bidwords that optimize both relevance and value, while adapting to real-time changes in their monetary value. Compared to ”SFT+Trie” (Exp.8), the ”SFT+WeightedTrie” configuration (Exp.10) achieves a 5.1 percentage point increase in hitrate@500 and substantially improves eCPM while concurrently maintaining rewrite relevance. Similarly, when employed in the generation phase alongside other generative rewriting methods (Exp.4 vs Exp.5, Exp.9 vs Exp.11), Weighted Trie significantly enhances key evaluation metrics associated with bidwords value. These results demonstrate that our Weighted Trie serves as an efficient, lightweight, and pluggable module that effectively enhances the value awareness of generative query rewriting methods. Notably, the incorporation of Weighted Trie significantly elevates the Spearman’s $\rho$, demonstrating that its value-relevance balancing mechanism during decoding enhances fine-grained value awareness.
4.4.3. Ablation Study of WDPO.
Traditional preference alignment techniques employed in BEQUE, such as DPO and Preference Ranking Optimization (PRO), are susceptible to significant overfitting during training, degrading performance in terms of both hitrate and relevance. The overfitting issue inherent to DPO arises from its tendency to overfit specific preference pairs in the training data, leading to compromised generalization capabilities. Consistent with this observation, Table 1 demonstrates marked declines in both hitrate and relevance when integrating DPO with either the ”SFT+Trie” (Exp.8 vs Exp.9) or the ”SFT+WeightedTrie” (Exp.10 vs Exp.11). In contrast, WDPO enables LLMs to focus learning efforts on high-value bidwords by attenuating the influence of preference pairs exhibiting minimal value differences during training. Consequently, the VALUE framework incorporating WDPO (Exp.12) achieves substantial gains across multiple value-centric metrics while preserving relevance.
@@
Figure 3. Trade-off between RPM and PV rele through $\Theta$. The specific values of RPM and PV rele are confidential, so the RPM and PV rele here are equivalently scaled.:
@@
$\Theta_{1}$: [0, 0, 0, 0, 0, 0, 0, …] 86.38% 75.31%
$\Theta_{2}$: [1, 1, 1, 1, 1, 1, 1, …] 89.65% 75.04%
$\Theta_{3}$: [1, 2, 3, 4, 5, 6, 7, …] 91.52% 74.55%
$\Theta_{4}$: [1, 2, 4, 8, 16, 32, 64, …] 88.63% 69.85%
$\Theta_{5}$: [2, 4, 8, 16, 32, 64, 128, …] 73.71% 60.53%